### v1

# Zero-Knowledge Proofs in Sublinear Space

Abstract
Modern zero-knowledge proof (ZKP) systems, essential for privacy and verifiable computation, suffer from a fundamental limitation: the prover typically uses memory that scales linearly with the computation’s trace length $T$, making them impractical for resource-constrained devices and prohibitively expensive for large-scale tasks. This paper overcomes this barrier by constructing, to our knowledge, the first sublinear-space ZKP prover. The core contribution is using an equivalence that reframes the algebraic process of generating a proof as an instance of the classic Tree Evaluation problem. Leveraging a recent space-efficient tree-evaluation algorithm, we design a streaming prover that recursively assembles the proof without ever materializing the full execution trace. The approach reduces prover memory from $\Theta(T)$ to $O(\sqrt{T})$ (up to lower-order polylogarithmic terms) while preserving identical proof size, verifier time, and transcript/security guarantees of the underlying system. This suggests transformation of verifiable computation from a specialized, server-bound task into one feasible on everyday devices, unlocking new applications in decentralized systems, on-device machine learning, and privacy-preserving technologies.
1 Introduction
Zero-knowledge proof systems, a cornerstone of modern cryptography, enable a prover to convince a verifier of a statement’s truth without revealing anything beyond that validity. Recent constructions—such as SNARKs and STARKs [1, 2]—achieve succinct proofs and fast verification, catalyzing applications from private cryptocurrency transactions [3] to verifiable machine learning and blockchain scalability.
Despite this progress, a fundamental bottleneck persists across practical systems: proof generation requires memory linear in the computation’s trace length. To prove a computation of trace length $T$, existing provers materialize the complete execution trace—i.e., machine state at every row in the algebraic intermediate representation (AIR)—consuming $\Theta(T)$ memory. This is not merely an implementation artifact: the algebraic machinery encodes this full trace into polynomials and enforces global relations (e.g., transition and permutation constraints). The memory cost limits provable scale, blocks deployment on mobile/embedded devices, and inflates costs for web-scale computations.
1.1 Prior Work and Technical Barriers
Current ZKPs are fundamentally trace-bound: their key operations require access to the entire trace.
  • Polynomial interpolation/FFT: Converting trace rows to evaluations or coefficients is typically implemented assuming random access to all $T$ entries; standard pipelines materialize the full trace in memory during these transforms.
  • Permutation arguments: Running-product accumulators enforce memory consistency via global products across all $T$ rows, again tying the prover to the whole trace.
  • Global checks: Boundary conditions and cross-block consistency induce access patterns that are difficult to implement in a strictly streaming fashion without redesign.
Prior attempts to reduce memory have improved constants or parallelization, but have not broken the $\Theta(T)$ barrier. We overcome this limitation by replacing global trace materialization with local, recursive cryptographic operations.
1.2 Main Contribution: A Sublinear-Space Prover
We establish a precise bridge between complexity theory and applied cryptography. We show that generating ZKP commitments is mathematically equivalent …Applying a recent space-efficient Tree Evaluation algorithm [4] yields, to our knowledge, the first sublinear-space prover within mainstream PCS-based SNARK/Plonkish arithmetizations that preserves verifier cost and (for linear PCSs) the proof/transcript of the baseline prover.
Theorem 1 (Main Result, Informal).
There exists a zero-knowledge proof system in which, for a computation of trace length $T$, the prover operates in $O(\sqrt{T}\cdot\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor). For linear PCS instantiations (e.g., KZG/IPA), the proof size and verifier complexity are unchanged and the transcript has the same distribution as in the standard prover (and is bit-for-bit identical in non-hiding variants).
Theorem 2 (Main Result, Formal).
Given a ZKP for arithmetic circuits instantiated with a PCS whose commit map is linear and computationally binding and a zero-knowledge instantiation (via a hiding PCS or standard masking), for any circuit of size/trace length $T$ we construct a prover $\mathcal{P}^{\prime}$ such that:
  1. 1.
    $\mathcal{P}^{\prime}$ uses $O(\sqrt{T}\cdot\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor).
  2. 2.
    The distribution of proofs $\pi^{\prime}$ generated by $\mathcal{P}^{\prime}$ is identical to that of the baseline prover $\mathcal{P}$ on the same public randomness (or programmed random oracle): in non-hiding PCSs commitments are deterministic and match bit-for-bit; in hiding PCSs the per-register blinder used by $\mathcal{P}^{\prime}$ is the sum of independent uniform block blinders and hence uniform, yielding identical commitment and opening distributions.
  3. 3.
    Soundness and zero-knowledge of the underlying protocol are preserved.
This represents a quadratic improvement in prover space, transforming terabyte-scale memory requirements into megabytes and making large-scale verifiable computation practical.
1.3 Technical Overview
This approach synthesizes ideas from complexity theory and cryptography in three steps:
  1. 1.
    Structural decomposition. Following Williams [5], we linearize the circuit into an AIR trace and impose a block-respecting structure with bounded cross-register read-degree and $k=O(1)$ registers (typical fixed-column AIRs), yielding a computation graph with bounded indegree.
  2. 2.
    Tree Evaluation Equivalence. We prove that the prover’s commitment generation factors into a recursive function over an implicit computation tree: each node consumes children’s commitments/auxiliary data and outputs a local block commitment (Section 4). Under a linear PCS, the resulting vector of trace-commitments has the same distribution as the baseline prover (and is bit-for-bit identical in non-hiding variants).
  3. 3.
    Space-efficient evaluation. We evaluate the implicit tree using the Cook–Mertz algorithm [4]. Balancing block size $b_{\text{blk}}$ against the $\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|$ stack term yields overall space $O\!\big{(}b_{\text{blk}}+\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|\big{)}$ for the traversal plus $O(b_{\text{blk}})$ for local work, optimized at $b_{\text{blk}}=\Theta(\sqrt{T})$, giving $O(\sqrt{T})$ space up to a $\log\log$ factor (Section 5). The prover time is $O(T\log N)\cdot\mathrm{poly}(\lambda)$ for evaluation-basis KZG/IPA with domain size $N\!\geq\!T$ (and $N=\mathrm{poly}(T)$ in typical settings).
1.4 Significance
This work establishes a bridge between the space complexity of abstract algorithms and that of concrete cryptographic protocols. It suggests that when the core relations of a protocol can be expressed as a function compatible with a space-efficient evaluation scheme for a computational model, the protocol can inherit that scheme’s space complexity. The quadratic reduction in prover memory enables new deployment scenarios—on-device proving, broader decentralization, and verifiable large-scale computation—impacting the economics and feasibility of ensuring computational integrity and privacy.
2 Preliminaries
In this section, we define the core concepts and establish the notation used throughout the paper. We model computations as arithmetic circuits, formalize the properties of polynomial commitment schemes, and define the Tree Evaluation problem.
2.1 Notation
We denote the security parameter by $\lambda$. Let $\mathbb{F}$ be a finite field with $|\mathbb{F}|\geq\mathrm{poly}(T)$ so that Schwartz–Zippel error is at most $1/\mathrm{poly}(T)$. Cryptographic parameters are chosen so that the commitment group $\mathcal{C}$ (used by the PCS below) has order at least $2^{\lambda}$. When convenient, we assume $\lambda\geq c\log T$, which implies $1/\mathrm{poly}(T)=2^{-\Omega(\lambda)}$; this relationship is not required but simplifies asymptotic notation.
If $T\mid(|\mathbb{F}|{-}1)$ we fix a primitive $T$-th root of unity $\omega\in\mathbb{F}$ and use the domain $H=\{\omega^{0},\ldots,\omega^{T-1}\}$. Otherwise we choose a domain of size $N\geq T$ either by (i) passing to an extension field $\mathbb{K}\supseteq\mathbb{F}$ that contains a multiplicative subgroup of order $N$, or (ii) using a size-$N$ coset domain inside $\mathbb{F}$ (typically $N$ a power of two), and we pad the trace to length $N$. All polynomial interpolations below are then taken over the chosen domain of size $N$, with degree <N<N; when $N=T$ this specializes to the first case.
All polynomials we manipulate have degree at most $D<N$, yielding Schwartz–Zippel error $D/|\mathbb{F}|\leq 1/\mathrm{poly}(T)$. A function $\nu:\mathbb{N}\to\mathbb{R}$ is negligible if for every positive polynomial $p(\cdot)$, there exists an integer $n_{0}$ such that for all $n>n_{0}$, $\nu(n)<1/p(n)$.
We use $t$ to denote the total number of steps of a Turing machine computation and $T$ for the size of the linearized circuit (equivalently, the number of rows in the execution trace). We use $k$ for the number of memory regions or registers. We use $\mathrm{poly}(\lambda)$ to denote an unspecified function that is polynomial in $\lambda$. We use $b_{\text{blk}}$ to denote computation block size and $b_{\text{val}}$ to denote the bit-length of values in tree evaluation.
2.2 Computational Model: Arithmetic Circuits
We model computations as arithmetic circuits over a finite field $\mathbb{F}$. This model is general and directly corresponds to the relations proven by many modern ZKP systems. We first establish the bridge between Turing machine computations and arithmetic circuits.
Lemma 1 (Turing Machine to Arithmetic Circuit Compilation).
Any $t$-step computation on a single-tape Turing machine can be efficiently compiled into an arithmetic circuit of size $T=O(t\log t)$ over a finite field $\mathbb{F}$. This transformation preserves the computational semantics and allows the structural decomposition techniques of Williams [5] to be applied to the circuit model.
Proof Sketch.
This follows from standard oblivious-simulation constructions: encode the time-$t$ TM run as a tableau and select the active cell/head position via $O(\log t)$-bit index decoding per step. Each local consistency check (adjacent rows/columns) is compiled into $O(1)$ arithmetic constraints over $\mathbb{F}$. The resulting circuit has size $T=O(t\log t)$ and evaluates gate-by-gate to reproduce the original computation. (With constant-tape or RAM models under oblivious access, one can obtain $T=O(t)$.) ∎
Definition 1 (Arithmetic Circuit).
An arithmetic circuit $C:\mathbb{F}^{n+n^{\prime}}\to\mathbb{F}^{m}$ is a directed acyclic graph where nodes are called gates. Gates have an indegree of 0 (input gates or constant gates), 1 (unary operations), or 2 (binary operations). Input gates are labeled with input variables from a public statement $x\in\mathbb{F}^{n}$ and a private witness $w\in\mathbb{F}^{n^{\prime}}$. Internal gates are labeled with arithmetic operations from $\{+,\times,-\}$ over $\mathbb{F}$; wires may fan out. A subset of gates are designated as output gates. The size of the circuit, after fixing a topological order, is denoted $T$ and equals the total number of gates.
Definition 2 (Execution Trace).
For an arithmetic circuit $C$ of size $T$ and a given assignment to its input gates, the execution trace $\tau$ is a matrix $W\in\mathbb{F}^{T\times k}$ with $k$ columns (registers) and $T$ rows (computation steps) indexed by $i\in\{0,\ldots,T{-}1\}$. Entry $W[i,j]$ is the value in register $j$ at step $i$, following the Algebraic Intermediate Representation (AIR). The correctness of a computation is captured by a fixed set of transition constraints: there exist polynomials $P_{1},\ldots,P_{r}$ such that for each $i\in\{0,\ldots,T{-}2\}$ and each $\ell\in[r]$,
$P_{\ell}\big{(}W[i,\cdot],W[i{+}1,\cdot]\big{)}=0,$
together with initialization and terminal constraints on $W[0,\cdot]$ and $W[T{-}1,\cdot]$. A valid trace $\tau$ satisfies all these constraints.
In standard ZKP constructions, the execution trace $\tau$ is encoded as a collection of trace polynomials, one per register/column. Specifically, fix a domain of size $N\geq T$ as described in the notation section. For each $j\in[k]$, define $F_{j}(X)$ to be the unique polynomial of degree <N<N that interpolates the $j$-th column over the domain $H=\{\omega^{0},\ldots,\omega^{N-1}\}$ (or the chosen coset domain), i.e., $F_{j}(\omega^{i})=W[i,j]$ for $i\in\{0,\ldots,T-1\}$ and equals the padded values for $i\in\{T,\ldots,N-1\}$. The prover’s main task is to commit to these trace polynomials.
2.3 Polynomial Commitment Schemes
We recall the standard interface and security properties for polynomial commitment schemes (PCSs) and fix the algebraic conventions we rely on throughout.
Definition 3 (Polynomial Commitment Scheme).
A PCS over a field $\mathbb{F}$ and maximum degree $D$ is a tuple of algorithms $(\mathsf{Setup},\mathsf{Commit},\mathsf{Open},\mathsf{Verify})$:
  • $\mathsf{Setup}(1^{\lambda},D)\to pp$: randomized, outputs public parameters $pp$.
  • $\mathsf{Commit}(pp,f;r)\to\mathsf{com}$: (possibly randomized) commitment to $f\in\mathbb{F}[X]$ with $\deg f\leq D$ using randomness $r$.
  • $\mathsf{Open}(pp,f,z;r)\to\pi$: opening proof for $y=f(z)$, using the same $r$ if needed.
  • $\mathsf{Verify}(pp,\mathsf{com},z,y,\pi)\in\{0,1\}$: verifies the opening.
Correctness.
For all valid inputs, if $\mathsf{com}\!\leftarrow\!\mathsf{Commit}(pp,f;r)$ and $\pi\!\leftarrow\!\mathsf{Open}(pp,f,z;r)$ with $y=f(z)$, then $\mathsf{Verify}(pp,\mathsf{com},z,y,\pi)=1$.
Definition 4 (Binding).
The binding game $\mathsf{Game}_{\mathcal{A}}^{\mathrm{Bind}}(1^{\lambda})$ proceeds as usual: $\mathcal{A}$ receives $pp\!\leftarrow\!\mathsf{Setup}(1^{\lambda},D)$ and wins if it outputs $(\mathsf{com},z,y,\pi,y^{\prime},\pi^{\prime})$ with $y\neq y^{\prime}$ but $\mathsf{Verify}(pp,\mathsf{com},z,y,\pi)=\mathsf{Verify}(pp,\mathsf{com},z,y^{\prime},\pi^{\prime})=1$. The scheme is computationally binding if every PPT $\mathcal{A}$ wins with probability negligible in $\lambda$.
Definition 5 (Hiding).
In the hiding game $\mathsf{Game}_{\mathcal{A}}^{\mathrm{Hide}}(1^{\lambda})$ the challenger commits to one of $\{f_{0},f_{1}\}$ chosen uniformly and $\mathcal{A}$ must guess the bit; the scheme is computationally hiding if $|\Pr[b{=}b^{\prime}]-\tfrac{1}{2}|$ is negligible.
Basis and SRS discipline.
Throughout we fix (and match the baseline prover’s) structured reference string (SRS) and the polynomial basis in which commitments are formed—either coefficient basis or an evaluation/Lagrange basis over a public domain $H=\{\omega^{0},\ldots,\omega^{N-1}\}$. When we commit in evaluation basis, we view a polynomial $f$ as its evaluation vector $(f(\omega^{i}))_{i\in[N]}$; when we commit in coefficient basis, we view $f$ by its coefficients. All linear relations claimed below are taken with respect to this fixed basis.
Linearity in message and randomness.
We assume the PCS commit map is linear in the committed message and (if hiding) in its randomness:
$\mathsf{Commit}(pp,f;r_{f})+\mathsf{Commit}(pp,g;r_{g})=\ \mathsf{Commit}\!\left(pp,\ f{+}g;\ r_{f}{+}r_{g}\right).$ (1)
For non-hiding KZG, commitments are deterministic and (1) holds with $r_{f}=r_{g}=0$. For hiding KZG/IPA variants, $r_{f},r_{g}\!\leftarrow\!\mathbb{F}$ independent uniform implies $r_{f}{+}r_{g}$ is uniform, preserving commitment distributions under linear aggregation.
Lemma 2 (Linear aggregation of block commitments).
Let $F=\sum_{t=1}^{B}G_{t}$ be a decomposition into block polynomials (all represented in the fixed basis). If $\mathsf{com}_{t}=\mathsf{Commit}(pp,G_{t};r_{t})$, then
$\sum_{t=1}^{B}\mathsf{com}_{t}\ =\ \mathsf{Commit}\!\left(pp,\ \sum_{t=1}^{B}G_{t};\ \sum_{t=1}^{B}r_{t}\right).$
If the scheme is hiding and each $r_{t}$ is independent uniform, then $\sum_{t}r_{t}$ is uniform, so the RHS has the same distribution as a baseline commitment to $F$ with fresh randomness.
Proof.
Apply (1) repeatedly. Uniformity of the sum in a finite field follows from independence. ∎
Corollary 1 (Identical transcript for linear PCSs).
Suppose the baseline prover commits to each trace/derived polynomial $F_{j}$ in the same basis/SRS fixed above. If $F_{j}=\sum_{t}G_{t,j}$ is the block decomposition and the block commitments $\{\mathsf{Commit}(G_{t,j};r_{t,j})\}_{t}$ are aggregated as in Lemma 2, then for non-hiding PCSs the commitment to $F_{j}$ is bit-for-bit identical to the baseline; for hiding PCSs it has the same distribution.
Quotient construction (coefficient basis).
Let $R$ denote the randomized linear combination of constraint residual polynomials and let $Z_{H}$ be the vanishing polynomial of $H$. We define $Q$ by the polynomial identity
$R(X)\;=\;Z_{H}(X)\cdot Q(X),$
and compute $Q$ in coefficient basis via a streaming backward division by $Z_{H}$, emitting coefficient blocks and committing to them blockwise. By Lemma 2, these block commitments aggregate to the baseline commitment to $Q$ (for linear PCSs), and the approach avoids divide-by-zero at points of $H$.
Streaming openings at a point.
To open a committed polynomial at a challenge point $\zeta\!\notin\!H$, a Lagrange/evaluation-basis prover can compute $f(\zeta)$ in a single streaming pass using barycentric evaluation, maintaining $O(1)$ accumulators; coefficient-basis provers use a Horner-style pass over streamed coefficients. In both cases, no full polynomial materialization is required, preserving the space bounds.
2.4 The Tree Evaluation Problem
The space-efficient prover relies on an algorithm for the Tree Evaluation problem.
Definition 6 (Tree Evaluation Problem).
An instance of the Tree Evaluation problem is a rooted tree $\mathcal{T}$ where:
  • Each leaf node is labeled with a value $x\in\{0,1\}^{\leq b_{\text{val}}}$ (a bitstring of length at most $b_{\text{val}}$).
  • Each internal node $v$ has an ordered list of $d_{v}\leq\Delta$ children and is labeled with an efficiently computable function
    $f_{v}:\ \big{(}\{0,1\}^{\leq b_{\text{val}}}\big{)}^{d_{v}}\ \to\ \{0,1\}^{\leq b_{\text{val}}}.$
  • The value of an internal node $v$ is defined recursively as $f_{v}$ applied to the values of its children.
Let $|\mathcal{T}|$ denote the number of nodes and $h$ the height. The goal is to compute the value of the root node.
The naive depth-first traversal algorithm for this problem requires $O(h\cdot b_{\text{val}})$ space. However, a recent breakthrough by Cook and Mertz provides a substantially more space-efficient solution.
Theorem 3 (Cook–Mertz Algorithm [4]).
Let $\mathcal{T}$ be a rooted tree with $|\mathcal{T}|=n$ nodes. The Tree Evaluation problem on $\mathcal{T}$ can be solved using
$O\!\big{(}\log n\cdot\log\log n\big{)}$
workspace for traversal (RAM model with word size $\Theta(\log n)$). In our setting, where node values have bit-length $b_{\text{val}}$ and each node has degree at most $\Delta$, buffering children’s values adds $O(\Delta\cdot b_{\text{val}})$, so the total space is
$O\!\big{(}\Delta\cdot b_{\text{val}}\;+\;\log|\mathcal{T}|\cdot\log\log|\mathcal{T}|\big{)}.$
Corollary (instantiation). With $\Delta=O(1)$ and $b_{\text{val}}=\Theta(\lambda)$, the Cook–Mertz stack uses $O\!\big{(}\lambda+\log|\mathcal{T}_{C}|\log\log|\mathcal{T}_{C}|\big{)}$ space. Since $\log|\mathcal{T}_{C}|\leq O(T/b_{\text{blk}})$, this stack term is $O\!\big{(}(T/b_{\text{blk}})\cdot\log\log|\mathcal{T}_{C}|\big{)}=O(T/b_{\text{blk}})$ up to a $\log\log$ factor; with $b_{\text{blk}}=\Theta(\sqrt{T})$, the overall space remains $O(\sqrt{T})$ up to a $\log\log$ factor, as in Theorem 5.
Parameter instantiation for our setting.
In our construction the computation tree has maximum degree $\Delta=O(k)$ (constant in typical AIRs), node value size $b_{\text{val}}=\Theta(\lambda)$ (a small tuple of commitments and $O(1)$ field elements), and height $h=O(T/b_{\text{blk}})$. Since $\log|\mathcal{T}|\leq h\log\Delta=O(h)$, Theorem 3 gives stack space
$O\!\big{(}\Delta\cdot b_{\text{val}}+\log|\mathcal{T}|\cdot\log\log|\mathcal{T}|\big{)}\;=\;O\!\big{(}\lambda+(T/b_{\text{blk}})\cdot\log\log|\mathcal{T}|\big{)},$
i.e., $O(\lambda+T/b_{\text{blk}})$ up to a $\log\log$ factor, which we combine with $O(b_{\text{blk}})$ local space and optimize at $b_{\text{blk}}=\Theta(\sqrt{T})$.
3 From Computations to Graphs
The first step in construction is to impose a regular, block-based structure onto an arbitrary computation. This process, adapted from the complexity-theoretic work of Hopcroft, Paul, and Valiant [7] and Williams [5], allows us to reason about local data dependencies. We show how to transform any arithmetic circuit into an equivalent “block-respecting” form and then define its corresponding computation graph.
3.1 Block-Respecting Decomposition
We conceptually partition the execution trace matrix of an arithmetic circuit into discrete blocks using the AIR (Algebraic Intermediate Representation) model established in Section 2. For $n\in\mathbb{N}$, write $n]\coloneqq\{1,\ldots,n\}$.
Definition 7 (Block-Respecting Computation).
A computation corresponding to an arithmetic circuit $C$ of size $T$ with execution trace matrix $W\in\mathbb{F}^{T\times k}$ is $(b_{\text{blk}},B)$-block-respecting, for parameters $b_{\text{blk}}$ (block size) and $B=\lceil T/b_{\text{blk}}\rceil$ (number of blocks), if:
  1. 1.
    Time Blocks: The $T$ rows of $W$ are partitioned into $B$ consecutive time blocks $W^{(1)},W^{(2)},\ldots,W^{(B)}$, where each $W^{(t)}\in\mathbb{F}^{b_{\text{blk}}\times k}$ contains at most $b_{\text{blk}}$ consecutive rows of the trace matrix.
  2. 2.
    Memory Regions: The $k$ columns of $W$ represent $k$ memory regions (registers). In standard fixed-column AIRs, $k=O(1)$; all bounds below scale with $k$ in general.
  3. 3.
    Locality (first-order transition with bounded cross-register reads): The AIR is first-order in time: there is a fixed set of transition polynomials $P_{1},\ldots,P_{r}$ such that each interior row pair (i,i+1)(i,i{+}1) satisfies $P_{\ell}(W[i,\cdot],W[i{+}1,\cdot])=0$ for all $\ell\in[r]$. Moreover, for each target register $m$, computing its next-row value uses the previous-row value of $m$ and at most $r_{\text{reg}}$ additional registers from that previous row, where $r_{\text{reg}}$ is the cross-register read-degree. In fixed-column AIRs, $r_{\text{reg}}=O(1)$; in general, bounds may scale with $k$.
For arithmetic circuits, this decomposition is achieved by partitioning the topologically ordered execution trace matrix into consecutive row blocks.
Lemma 3 (Block-Respecting Transformation).
Any arithmetic circuit $C$ of size $T$ with execution trace matrix $W\in\mathbb{F}^{T\times k}$ naturally yields a $(b_{\text{blk}},\lceil T/b_{\text{blk}}\rceil)$-block-respecting computation with no asymptotic overhead in time or space.
Proof.
Consider the execution trace matrix $W\in\mathbb{F}^{T\times k}$ for circuit $C$:
  1. 1.
    Matrix Partitioning: Partition the $T$ rows into $B=\lceil T/b_{\text{blk}}\rceil$ consecutive blocks of size at most $b_{\text{blk}}$.
  2. 2.
    Block-Respecting Property: By the AIR structure (Section 2), the transition relation is first-order in time and, for each target register, depends on its own previous-row value and at most $r_{\text{reg}}$ additional previous-row registers (with $r_{\text{reg}}=O(1)$ in fixed-column AIRs). Thus, within each block, dependencies are confined to the block together with boundary values from the immediately preceding block.
  3. 3.
    Evaluation: The original evaluator processes the trace row-by-row; grouping rows into blocks preserves semantics and does not change time or space complexity up to constant factors.
3.2 The Computation Graph
From a block-respecting computation, we construct a graph that captures the flow of information between blocks using the AIR trace structure.
Definition 8 (Computation Graph).
Let $C$ be a $(b_{\text{blk}},B)$-block-respecting computation with execution trace matrix $W\in\mathbb{F}^{T\times k}$. The computation graph $G_{C}=(V,E)$ is a directed acyclic graph defined as follows:
  • Nodes $V$: The set of nodes $V$ contains two types of vertices:
    • Computation Nodes: A node $(m,t)$ for each register $m\in[k]$ and time block $t\in[B]$. This node represents the state of register $m$ within the $t$-th block of rows.
    • Source Nodes: A node $(m,0)$ for each register $m\in[k]$, representing the initial contents of register $m$ (public inputs or private witness values).
  • Edges $E$: An edge $(u,v)\in E$ exists if the computation at node $v$ directly requires data from node $u$. Formally, for a computation node $v=(m,t)$ with $t>0$:
    1. 1.
      There is an edge from (m,t1)(m,t-1) to $(m,t)$, representing the dependency on the same register from the previous time block.
    2. 2.
      For each register $m^{\prime}\neq m$ whose boundary value at the end of block t1t{-}1 is read when producing block $t$ for register $m$, there is an edge from $(m^{\prime},t-1)$ to $(m,t)$.
    3. 3.
      For $t=1$, there are edges from the appropriate source nodes $(m^{\prime},0)$ to $(m,1)$ as determined by the circuit’s input dependencies.
Crucially for the space-efficient prover, the structure of $G_{C}$ need not be stored. Given the static description of the arithmetic circuit $C$ and an identifier for a node $(m,t)$, its predecessors can be enumerated in $O(\log T)$ space by scanning the local transition template and the circuit’s wiring.
Lemma 4 (Bounded Indegree).
Each computation node $(m,t)$ in $G_{C}$ has indegree at most $\min\{k,\,1{+}r_{\text{reg}}\}$, where $r_{\text{reg}}$ is the cross-register read-degree from the definition above.
Proof.
By first-order locality with bounded cross-register reads, producing block $t$ for register $m$ depends on the boundary value of register $m$ from block t1t{-}1 and on at most $r_{\text{reg}}$ other registers’ boundary values from block t1t{-}1. Thus the number of incoming edges to $(m,t)$ is at most $1{+}r_{\text{reg}}$, and certainly no more than $k$. ∎
We define a function $\mathsf{SimBlock}$ that represents the local computation occurring at each node.
Algorithm 1 SimBlock
0: Boundary vectors $\{B_{0},B_{1},\ldots,B_{r_{\text{reg}}}\}$ from predecessor nodes in block t1t{-}1 (including $B_{0}$ from (m,t1)(m,t{-}1)), time block index $t$, register $m$
0: Output data block $D_{\text{out}}\in\mathbb{F}^{b_{\text{blk}}}$ for register $m$ in time block $t$
[…]
2:for each row $i$ in time block $t$ do
3:  $\text{state}\leftarrow\mathsf{F}\big{(}\text{state}\big{)}$ {apply the local transition map induced by the fixed polynomials $P_{1},\ldots,P_{r}$}
4:  $D_{\text{out}}[i]\leftarrow\text{state}[m]$
5:end for
6:return  $D_{\text{out}}$
Lemma 5 (Graph Simulability).
Given the contents of all predecessor nodes of a computation node $(m,t)$, the content of $(m,t)$ can be computed in time $O\!\big{(}(1{+}r_{\text{reg}})\,b_{\text{blk}}\big{)}$ and space $O\!\big{(}b_{\text{blk}}+\deg^{-}(m,t)\big{)}$ using Algorithm 1. In particular, if the AIR has constant read-degree $r_{\text{reg}}=O(1)$ (hence $\deg^{-}(m,t)=O(1)$), this is $O(b_{\text{blk}})$ time and $O(b_{\text{blk}})$ space.
Proof.
Algorithm 1 takes as input only the predecessor boundary vectors, whose total size is $O(\deg^{-}(m,t))$. Each iteration updates the local state via a constant-arity transition map $\mathsf{F}$ (dependent only on $P_{1},\ldots,P_{r}$) and writes one entry of $D_{\text{out}}$, yielding $b_{\text{blk}}$ iterations and the stated time bound. Storing the predecessor boundaries and the output requires $O(\deg^{-}(m,t)+b_{\text{blk}})$ field elements. ∎
4 The Tree Evaluation Equivalence
This section presents this work’s main conceptual contribution. We show that the process of generating a zero-knowledge proof can be reframed as an evaluation on an implicit tree derived from the computation graph. This reframing replaces the monolithic generation of the trace with a recursive, space-efficient construction of cryptographic commitments.
4.1 From Computation Graphs to Implicit Trees
The computation graph $G_{C}$ is a directed acyclic graph (DAG). We can “unroll” this DAG into an equivalent tree structure, where shared nodes in the DAG are duplicated.
Definition 9 (Computation Tree).
Given a computation graph $G_{C}$ and a designated final output node $v_{\mathrm{root}}\in G_{C}$ (e.g., $(1,B)$), the corresponding computation tree $\mathcal{T}_{C}$ is an implicitly defined tree where:
  • The root of $\mathcal{T}_{C}$ corresponds to $v_{\mathrm{root}}$ in $G_{C}$.
  • For any node $v$ in $\mathcal{T}_{C}$ corresponding to a node $u\in G_{C}$, its children in $\mathcal{T}_{C}$ correspond to the predecessors of $u$ in $G_{C}$.
  • The leaves of $\mathcal{T}_{C}$ correspond to the source nodes of $G_{C}$.
We never materialize $\mathcal{T}_{C}$ explicitly. The tree view is used to establish distributional equivalence of transcripts; for execution, we evaluate the underlying DAG layer-by-layer so that each node is computed exactly once and retained only until its $O(1)$ consumers read its $\mathsf{aux}$. This schedule achieves the same asymptotic stack bound as the tree traversal (up to an additive $O(\lambda)$ buffer). The height is $h=B=O(T/b_{\text{blk}})$ and the branching factor is $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}=O(k)$ (see §3).
4.2 Tree Functions as Cryptographic Generators
We lift the computation from field elements to cryptographic state. Each node of the computation tree returns (i) a one-hot vector of commitments (nonzero only at its own register coordinate) and (ii) a compact auxiliary state $\mathsf{aux}$ carrying boundary data and running accumulators for global checks (permutation, lookups).
Definition 10 (Auxiliary state).
For a node $v=(m,t)$ (register $m$, time-block $t$), let
$\mathsf{aux}_{v}\ =\ \big{(}\ \mathsf{boundary}_{v}\in\mathbb{F}^{k},\ \ Z_{v}\in\mathbb{F},\ \ Z^{\mathrm{L}}_{v}\in\mathbb{F}^{\leq c_{\mathrm{L}}}\ \big{)},$
where $\mathsf{boundary}_{v}$ stores the end-of-block values of all $k$ registers, $Z_{v}$ is the permutation product accumulator, and $Z^{\mathrm{L}}_{v}$ is an optional tuple of lookup accumulators as required by the chosen lookup scheme; $c_{\mathrm{L}}=O(1)$.
Block-aligned domain slices.
Fix a partition $H=\bigsqcup_{t=1}^{B}H_{t}$ into contiguous subdomains aligned with time blocks (padding the final slice if N>TN\!>\!T). For register $m$ and block $t$, let $G_{t,m}$ be the degree-<N<N polynomial that equals the $m$-th register’s values on $H_{t}$ and 0 elsewhere.
Algorithm 2 $\mathsf{EvalBlock}(m,t,\text{inputs})$
0: Boundary vectors from predecessors (including (m,t1)(m,t{-}1)), block index $t$, target register $m$
0:local (rowwise values needed for checks), $\mathsf{boundary}_{\text{out}}$, and the length-$|H_{t}|$ vector for register $m$
1: Assemble prior-row inputs at the block boundary
2:for each row $i\in H_{t}$ in order do
3:  Advance the AIR state via the fixed polynomials $P_{1},\ldots,P_{r}$
4:  Emit the row-local fields required by global checks into local; emit the $m$-coordinate into the register-$m$ output
5:end for
6:return  $(\text{local},\ \mathsf{boundary}_{\text{out}},\ \text{regM\_vals})$
Permutation/lookup updates are multiplicative and block-factored.
We formalize the streaming of global accumulators. For concreteness, consider a Plonk-style permutation with registers indexed by $c\in[k]$, a column id mapping $\mathrm{id}_{c}(i)$ and a permutation $\sigma_{c}(i)$. Let $w_{c}(i)$ be the value in register $c$ at row $i$.
Lemma 6 (Blockwise permutation update).
Fix $(\beta,\gamma)\in\mathbb{F}^{2}$ and a block $H_{t}=\{i_{t},\ldots,i_{t}+|H_{t}|-1\}$. The permutation accumulator $Z$ satisfies
$Z(i{+}1)\ =\ Z(i)\cdot\frac{\prod_{c=1}^{k}\big{(}w_{c}(i)+\beta\cdot\mathrm{id}_{c}(i)+\gamma\big{)}}{\prod_{c=1}^{k}\big{(}w_{c}(i)+\beta\cdot\sigma_{c}(i)+\gamma\big{)}}\qquad(i\in H_{t}),$
and thus admits a multiplicative block factor
$Z_{\mathrm{end}}^{(t)}\ =\ Z_{\mathrm{start}}^{(t)}\cdot F_{t}(\beta,\gamma),\qquad F_{t}(\beta,\gamma)=\prod_{i\in H_{t}}\frac{\prod_{c}\big{(}w_{c}(i)+\beta\,\mathrm{id}_{c}(i)+\gamma\big{)}}{\prod_{c}\big{(}w_{c}(i)+\beta\,\sigma_{c}(i)+\gamma\big{)}}.$
Consequently, concatenating blocks in time order yields the same $Z$ as the monolithic run; commutativity of multiplication implies the same final value for any block schedule consistent with row order.
Lemma 7 (Blockwise lookup update).
For a standard lookup accumulator $Z_{\mathrm{L}}$ (e.g., Plookup/Halo2-style), there exists a per-row multiplicand $\phi^{\mathrm{L}}(i)$ determined by the scheme’s lookup relation such that
$Z_{\mathrm{L}}(i{+}1)\ =\ Z_{\mathrm{L}}(i)\cdot\phi^{\mathrm{L}}(i)\qquad(i\in H_{t}),$
whence $Z_{\mathrm{L,end}}^{(t)}=Z_{\mathrm{L,start}}^{(t)}\cdot\prod_{i\in H_{t}}\phi^{\mathrm{L}}(i)$. Thus lookup accumulators also stream with multiplicative block factors.
Quotient construction (coefficient basis, no divide-by-zero).
Let $R$ denote the standard randomized linear combination of constraint residual polynomials and $Z_{H}$ the vanishing polynomial of $H$. The quotient $Q$ is defined by the polynomial identity $R=Z_{H}\cdot Q$ (no pointwise division on $H$). When $H$ is a multiplicative subgroup or coset, $Z_{H}(X)=X^{N}-c$ is monic of degree $N$, so dividing $R$ by $Z_{H}$ reduces to a linear-time backward sweep on coefficients. We therefore commit to $Q$ in coefficient basis (matching the baseline prover), and compute it in $O(\sqrt{T})$ space via a blocked sweep that emits coefficient blocks and aggregates their commitments using Lemma 2. This avoids division-by-zero at evaluation points and preserves the baseline commitment and transcript for linear PCSs.
Algorithm 3 Commitment tree function $F_{v}$ at node $v=(m,t)$
0: For each child $c_{i}$: $(\boldsymbol{\mathsf{com}}_{i},\mathsf{aux}_{i})$ with $\boldsymbol{\mathsf{com}}_{i}\in\mathcal{C}^{k}$ one-hot
[…]
3:Permutation/lookup checks: using local, update
$Z_{\text{out}}\leftarrow Z_{\text{in}}\cdot F_{t}(\beta,\gamma),\qquad Z^{\mathrm{L}}_{\text{out}}\leftarrow Z^{\mathrm{L}}_{\text{in}}\cdot\prod_{i\in H_{t}}\phi^{\mathrm{L}}(i),$
and abort if any consistency check fails
4: Form $G_{t,m}$ by interpolating regM_vals on $H_{t}$ and padding 0 off $H_{t}$
5: Sample $r_{v}\leftarrow\mathbb{F}$ if hiding PCS, else $r_{v}\leftarrow 0$
[…]
7: Initialize $\boldsymbol{\mathsf{com}}_{\mathrm{out}}\leftarrow\boldsymbol{0}\in\mathcal{C}^{k}$
8: Let $c^{\star}$ be the child with the same register $m$ (if $t>1$), else $c^{\star}\leftarrow\bot$
[…]
11:return  $(\boldsymbol{\mathsf{com}}_{\mathrm{out}},\mathsf{aux}_{\mathrm{out}})$
Aggregation discipline and final aggregator.
At each $(m,t)$ we carry forward only coordinate $m$ and never duplicate commitments across registers. A virtual aggregator node $v_{\mathrm{agg}}$ with children $\{(m,B)\}_{m\in[k]}$ reassembles the final $k$-tuple by placing each incoming $m$-coordinate into position $m$. By Lemma 2 and Corollary 1, for linear PCSs the resulting tuple matches (in distribution, and bit-for-bit for non-hiding) the baseline commitments to the global polynomials.
4.3 The Main Equivalence Theorem
Theorem 4 (Tree Evaluation Equivalence).
Let $C$ be a computation with execution trace length $T$ over a domain $H=\{\omega^{0},\dots,\omega^{N-1}\}$ of size $N\geq T$ as in Section 2, with trace polynomials $\{F_{j}(X)\}_{j=1}^{k}$ of degree <N<N. There exists a computation tree $\mathcal{T}_{C}$ and functions $\{F_{v}\}$ such that:
  1. 1.
    Evaluating $\mathcal{T}_{C}$ with a final aggregator node $v_{\mathrm{agg}}$ (whose children are $\{(m,B)\}_{m\in[k]}$) yields a $k$-tuple $\boldsymbol{\mathsf{com}}_{\mathrm{root}}$ whose distribution is identical to $(\mathsf{Commit}(pp,F_{1};R_{1}),\ldots,\mathsf{Commit}(pp,F_{k};R_{k}))$, where the $R_{j}$ follow the PCS randomness (uniform in hiding variants; $R_{j}\!=\!0$ in non-hiding variants).
  2. 2.
    $\mathcal{T}_{C}$ has height $h=O(T/b_{\text{blk}})$ and maximum degree $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}=O(k)$. Each $F_{v}$ is computable in $O\!\big{(}b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big{)}$ time and $O(b_{\text{blk}})$ space, where $C_{\text{commit}}(b)=O(\mathrm{MSM}(b))=O(b)$ with a preprocessed evaluation-basis SRS (conservatively, $O(b\cdot\mathrm{polylog}\,N)$).
Proof.
Parameters. Height and degree follow from $G_{C}$ (Section 3): $h=B=O(T/b_{\text{blk}})$ and $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}$. By Lemma 5, evaluating a block uses $O(b_{\text{blk}})$ space and $O((1{+}r_{\text{reg}})b_{\text{blk}})=O(b_{\text{blk}})$ time. Commitment time is $C_{\text{commit}}(b_{\text{blk}})$ as stated.
Block polynomials and decomposition. Partition $H$ into contiguous subdomains $H_{1},\ldots,H_{B}$ with $|H_{i}|\in\{b_{\text{blk}},\,N-(B-1)b_{\text{blk}}\}$ (padding if $N>T$). For each block $i$ and register $j$, define $G_{i,j}$ to be the unique degree <N<N polynomial with evaluations
$G_{i,j}(\omega^{\ell})=\begin{cases}W[\ell,j],&\omega^{\ell}\in H_{i},\\ 0,&\omega^{\ell}\notin H_{i}.\end{cases}$
By construction, $\sum_{i=1}^{B}G_{i,j}$ and $F_{j}$ agree on all $\omega^{\ell}\in H$.
Lemma 8 (Uniqueness over $H$).
If $p,q$ have degree <N<N and $p(\omega^{\ell})=q(\omega^{\ell})$ for all $\omega^{\ell}\in H$ with $|H|=N$, then $p\equiv q$.
Applying Lemma 8 gives $F_{j}(X)=\sum_{i=1}^{B}G_{i,j}(X)$ as polynomials.
Induction over $\mathcal{T}_{C}$ (commitments). Leaves (source nodes) output $(\boldsymbol{0},\mathsf{aux})$, i.e., the all-zero commitment vector together with auxiliary data encoding initial boundary values and the starting accumulator. For an internal node $v=(m,t)$ with same-register predecessor $c^{\star}=(m,t{-}1)$ when $t>1$, Algorithm 3 computes $\mathsf{com}_{t,m}=\mathsf{Commit}(G_{t,m};r_{v})$ and sets
$\boldsymbol{\mathsf{com}}_{\text{out}}[m]\;=\;\big{(}\boldsymbol{\mathsf{com}}_{c^{\star}}[m]\text{ if }t>1\text{ else }0_{\mathcal{C}}\big{)}\;+\;\mathsf{com}_{t,m},\quad\boldsymbol{\mathsf{com}}_{\text{out}}[\ell]=0\ \ (\ell\neq m).$
Thus the $m$-coordinate at $(m,t)$ commits to $\sum_{i\leq t}G_{i,m}$. The aggregator $v_{\mathrm{agg}}$ collects the $k$ coordinates from $\{(m,B)\}$ to form $\big{(}\mathsf{Commit}(F_{1};\cdot),\ldots,\mathsf{Commit}(F_{k};\cdot)\big{)}$; in hiding PCSs, per-coordinate blinders sum to a uniform blinder; in non-hiding PCSs commitments are deterministic in both constructions.
Induction over $\mathcal{T}_{C}$ (consistency). The local $\mathsf{ConsistencyCheck}$ enforces boundary matching and the correct multiplicative update of the permutation accumulator by the block factor derived from local. Strong induction from leaves upward shows that if all local checks pass, the concatenated trace satisfies the global AIR and permutation constraints; conversely, root consistency forces every child’s $\mathsf{aux}$ to agree with the unique global trace, preventing double-counting across duplicated subtrees.
Combining the two inductions completes the proof. ∎
5 The Sublinear-Space Prover
By combining the Tree Evaluation Equivalence (Theorem 4) with the space-efficient Cook–Mertz algorithm (Theorem 3), we construct a sublinear-space zero-knowledge prover. The core idea is to execute the proof generation not by materializing the trace, but by evaluating a implicit cryptographic computation tree.
5.1 The Streaming Prover Algorithm
We now give a phase-accurate, Fiat–Shamir–compatible schedule that evaluates the implicit tree in sublinear space and preserves the baseline transcript (for linear PCSs). The prover never materializes the full trace, instead making a small number of streaming passes whose memory footprint is $O(\sqrt{T})$ up to a $\log\log$ factor.
Parameters.
Let $b_{\text{blk}}=\lfloor\sqrt{T}\rfloor$, $B=\lceil T/b_{\text{blk}}\rceil$, and fix $\lambda=\Theta(\log T)$. We use the block-respecting graph $G_{C}$ and the implicit computation tree $\mathcal{T}_{C}$ rooted at the final node (Section 3).
Fiat–Shamir streaming schedule (Plonkish outline).
We mirror the baseline order of commitments and challenges:
  1. 1.
    Phase A (Selector/public commits). Commit once to public polynomials (selectors, fixed tables); no trace needed.
  2. 2.
    Phase B (Wire commits, no challenges). For each register $m$ and block $t$, compute $G_{t,m}$ via $\mathsf{EvalBlock}$ (Alg. 2); commit and add into the running $m$-coordinate using $F_{v}$ (Alg. 3). By Lemma 2, the aggregated commitment equals (in distribution) the baseline wire commitment.
  3. 3.
    Phase C (Permutation after sampling $(\beta,\gamma)$). Sample $(\beta,\gamma)$ via FS. Re-stream each block to compute the permutation block factor $F_{t}(\beta,\gamma)$ (Lemma 6); multiply into the running accumulator $Z$ and commit to $Z$ blockwise, aggregating linearly to the global $Z$ commitment.
  4. 4.
    Phase D (Quotient after sampling $\alpha$). Sample $\alpha$. Compute and commit to $Q$ in coefficient basis by streaming division of $R$ by $Z_{H}$: perform a blocked backward sweep over the coefficients of $R$, emitting coefficient blocks of $Q$ and aggregating their commitments linearly. No materialization of the full coefficient vector is required; the working set is $O(b_{\text{blk}})$.
  5. 5.
    Phase E (Openings after sampling $\zeta,\ldots$). Sample evaluation points $(\zeta,\ldots)$. For each polynomial that must be opened at $\zeta$, compute $f(\zeta)$ in a single streaming pass (barycentric in Lagrange basis; Horner in coefficient basis) without storing $f$. Produce the standard batched opening proofs.
Because the commitments in Phases A–D match the baseline (Cor. 1), the Fiat–Shamir challenges have identical distribution, and thus the final transcript distribution is preserved.
Algorithm 4 Sublinear-Space ZKP Prover (FS-ordered streaming)
0: Arithmetic circuit $C$ of trace length $T$, witness $w$, public params pppp
0: Proof $\pi$
1:Params: $b_{\text{blk}}:=\lfloor\sqrt{T}\rfloor$, $B:=\lceil T/b_{\text{blk}}\rceil$; define $G_{C}$ and the oracles for $\mathcal{T}_{C}$
2:Phase A: Commit to public polynomials (selectors, tables)
3:Phase B (wires): For $t=1$ to $B$ and each register $m\in[k]$: compute $(\text{local},\mathsf{boundary}_{\text{out}},\text{regM\_vals})\leftarrow\mathsf{EvalBlock}(m,t,\cdot)$; form $G_{t,m}$; $\mathsf{com}_{t,m}\leftarrow\mathsf{Commit}(pp,G_{t,m};r_{t,m})$; aggregate into the $m$-coordinate using $F_{v}$
4: Sample $(\beta,\gamma)$ via FS
5:Phase C (permutation): For $t=1$ to $B$: compute $F_{t}(\beta,\gamma)$ by re-streaming block $t$; update $Z\leftarrow Z\cdot F_{t}$; commit to $Z$ blockwise and aggregate linearly
6: Sample $\alpha$ via FS
7:Phase D (quotient): Stream the coefficients of $R$ and perform a blocked backward division by $Z_{H}$ to emit coefficient blocks of $Q$; commit to each block in coefficient basis and aggregate linearly
8: Sample $(\zeta,\ldots)$ via FS
9:Phase E (openings): For each required $f$: compute $f(\zeta)$ in a single streaming pass; produce batched openings $\pi_{\mathrm{open}}$
10: Assemble $\pi$ from (aggregated) commitments, evaluation claims, and openings; return $\pi$
Space and time.
The live workspace at any point consists of $O(b_{\text{blk}})$ local state, $O(1)$ accumulators for permutation/lookup factors, and the Cook–Mertz traversal stack of size $O(\log|\mathcal{T}_{C}|\log\log|\mathcal{T}_{C}|)=O(T/b_{\text{blk}})$ up to a $\log\log$ factor (constant arity). Hence total space
$S(T)\ =\ O\!\big{(}b_{\text{blk}}+\tfrac{T}{b_{\text{blk}}}+\lambda\big{)}\ =\ O(\sqrt{T})\ \text{up to a }\log\log\text{ factor}.$
With an evaluation-basis SRS, each block commit costs $O(b_{\text{blk}})$ MSMs; across $kB=\Theta(T/b_{\text{blk}})$ blocks this is $O(T)$ MSMs, plus linear-time streaming for permutation/quotient passes and standard batched openings.
5.2 Oracle Implementation Details
Each oracle is space-efficient and avoids storing the full computation graph:
  • $\mathcal{O}_{\mathrm{children}}(v)$: Represent $v$ by a path descriptor in $O(\log T)$ space. Predecessors are derived by scanning the transition template from the AIR and the circuit wiring, requiring only streaming access to $C$.
  • $\mathcal{O}_{\mathrm{leaf}}(v)$: For source nodes, read the $O(k)$ initial boundary values needed to seed block t=1t{=}1 and set the starting permutation accumulator (typically $1$). No commitments are created at leaves. Space: $O(k)=O(1)$.
  • $\mathcal{O}_{\mathrm{func}}(v,\cdot)$: Implements Algorithm 3; evaluates the block and performs interpolation/commitment using $O(b_{\text{blk}})$ space.
5.3 Complexity Analysis
Theorem 5 (Sublinear Space Complexity).
With $\lambda=\Theta(\log T)$, Algorithm 4 uses
$S(T)=O\!\big{(}\sqrt{T}\cdot\log\log T\ +\ \log T\big{)}=O\!\big{(}\sqrt{T}\cdot\log\log T\big{)}\ \text{ space},$
i.e., $O(\sqrt{T})$ up to a $\log\log$ factor.
Proof.
Total space is the sum of oracle space and the Cook–Mertz stack space.
  • Oracle space. $\mathcal{O}_{\mathrm{func}}$ uses $O(b_{\text{blk}})$ space (Algorithm 3 and Lemma 5); $\mathcal{O}_{\mathrm{children}}$ and $\mathcal{O}_{\mathrm{leaf}}$ add $O(\log T)$ and $O(1)$, respectively. Hence $S_{\mathrm{oracle}}=O(b_{\text{blk}})$.
  • Cook–Mertz space. By Theorem 3 (Section 2),
    $S_{\mathrm{CM}}=O\!\big{(}\Delta\cdot b_{\text{val}}+\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|\big{)}.$
    Here $\Delta=O(k)=O(1)$ and $b_{\text{val}}=\Theta(\lambda)$ (a small tuple of commitments and $O(1)$ field elements). The tree size satisfies $\log|\mathcal{T}_{C}|\leq h\cdot\log\Delta=O(h)$ with $h=O(T/b_{\text{blk}})$. Thus
    $S_{\mathrm{CM}}=O\!\big{(}\lambda+(T/b_{\text{blk}})\cdot\log\log|\mathcal{T}_{C}|\big{)}=O\!\big{(}\lambda+T/b_{\text{blk}}\big{)}\ \text{up to a }\log\log\text{ factor}.$
  • Total and optimization. Therefore
    $S(T)=O\!\big{(}b_{\text{blk}}+\lambda+T/b_{\text{blk}}\big{)}.$
    Choosing $b_{\text{blk}}=\lfloor\sqrt{T}\rfloor$ minimizes the first and third terms, giving
    $S(T)=O\!\big{(}\sqrt{T}+\lambda\big{)}=O\!\big{(}\sqrt{T}+\log T\big{)}=O(\sqrt{T}).$
Theorem 6 (Prover Time Complexity).
Algorithm 4 runs in time
$O\!\Big{(}\tfrac{T}{b_{\text{blk}}}\cdot\big{(}b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big{)}\Big{)}\cdot\mathrm{poly}(\lambda),$
where $C_{\text{commit}}(b)=O(\mathrm{MSM}(b))=O(b)$ with a preprocessed evaluation-basis SRS (conservatively, $O(b\cdot\mathrm{polylog}\,N)$). With $b_{\text{blk}}=\Theta(\sqrt{T})$ and $\lambda=\Theta(\log T)$, this yields $O\!\big{(}T\cdot\mathrm{polylog}\,N\big{)}$ time; the wire-commit portion is $O(T)$ MSMs.
Proof.
Each block/register pair $(t,m)$ triggers one call to $\mathcal{O}_{\mathrm{func}}$ (Algorithm 3), costing
$O(b_{\text{blk}})\ \text{(block evaluation)}\ +\ O\!\big{(}C_{\text{commit}}(b_{\text{blk}})\big{)}\ \text{(interpolation/commit)}.$
There are $kB=\Theta(T/b_{\text{blk}})$ such calls with $k=O(1)$, so the total field/group work is
$\Theta\!\Big{(}\frac{T}{b_{\text{blk}}}\cdot\big{(}b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big{)}\Big{)}.$
PCS operations contribute only scheme-dependent $\mathrm{poly}(\lambda)$ factors. Substituting $b_{\text{blk}}=\Theta(\sqrt{T})$ yields $O(T\log N)$ with the stated commitments.
Memoization schedule inside the evaluation.
Claim. There exists an evaluation schedule such that each node $(m,t)$ is computed exactly once and retained only until its $O(1)$ consumers at layer t+1t{+}1 have consumed its $\mathsf{aux}$; consequently at most a constant number of blocks per layer are live in addition to the Cook–Mertz stack.
Justification. Process the computation graph in increasing $t$; within each layer, any topological order suffices. By first-order locality with read-degree $r_{\text{reg}}=O(1)$, a node $(m,t)$ is needed only by its own successor (m,t+1)(m,t{+}1) and by at most $r_{\text{reg}}$ cross-register consumers at the same layer increment. Once those $O(1)$ consumers have read its $\mathsf{aux}$, $(m,t)$ can be discarded. This yields exactly $kB=\Theta(T/b_{\text{blk}})$ invocations of $\mathcal{O}_{\mathrm{func}}$ without duplication and adds only $O(1)$ blocks of live state beyond the $O(\lambda+T/b_{\text{blk}})$ stack space of Cook–Mertz, preserving the overall $O(\sqrt{T})$ space bound up to a $\log\log$ factor. ∎
6 Security Preservation
A critical aspect of our construction is that it achieves its space efficiency without compromising security. We prove that the sublinear-space prover inherits completeness, soundness, and zero-knowledge from the underlying PCS-IOP/NIZK protocol (instantiated over the AIR), and that the change of evaluation order (via tree evaluation) leaves transcript distributions unchanged (Theorem 4).
6.1 Completeness
Theorem 7 (Completeness Preservation).
If the underlying proof system is complete, then for any valid statement–witness pair $(x,w)\in R$, Algorithm 4 produces a proof $\pi$ that the verifier accepts with the same completeness guarantee as the underlying system (perfect completeness: acceptance with probability $1$; otherwise, the identical acceptance probability over the verifier’s randomness / Fiat–Shamir challenges).
Proof.
Completeness follows from correctness at each layer:
  1. 1.
    The block-respecting transformation (Lemma 3) is semantics-preserving.
  2. 2.
    The Tree Evaluation Equivalence (Theorem 4) shows that evaluating the computation tree yields the same $k$ trace-commitments as the standard linear-space prover, together with consistent auxiliary data.
  3. 3.
    The Cook–Mertz procedure faithfully evaluates the (implicit) tree to the same root values.
  4. 4.
    The final proof $\pi$ is therefore generated from the identical transcript distribution as in the standard prover.
By completeness of the underlying system, the verifier accepts with the same guarantee. ∎
6.2 Soundness
We model soundness against a prover that outputs an accepting proof for a false statement. The security parameter and degree bound are as in Section 2.
Definition 11 (Soundness Game).
The game $\mathsf{Game}_{\mathcal{A}}^{\mathrm{Sound}}(1^{\lambda},C)$:
  1. 1.
    The challenger samples public parameters $pp\leftarrow\mathsf{Setup}(1^{\lambda},D)$ (for a maximum polynomial degree $D<N$) and fixes an evaluation domain $H=\{\omega^{0},\ldots,\omega^{N-1}\}$ of size $N\geq T$.
  2. 2.
    The adversary $\mathcal{A}(pp)$ outputs a statement $x$ and purported proof $\pi$.
  3. 3.
    $\mathcal{A}$ wins if $\mathsf{Verify}(pp,x,\pi)=1$ and $x\notin L(R)$.
Theorem 8 (Soundness Preservation).
Let $\Pi_{\mathrm{std}}$ be the underlying PCS-IOP/NIZK instantiated for the AIR, and let $\Pi_{\mathrm{str}}$ be the streaming/tree-evaluation prover from Algorithm 4. If $\Pi_{\mathrm{std}}$ is computationally sound (under the binding of the PCS and its algebraic checks), then $\Pi_{\mathrm{str}}$ is computationally sound. In particular, any PPT adversary $\mathcal{A}$ that wins $\mathsf{Game}_{\mathcal{A}}^{\mathrm{Sound}}$ against $\Pi_{\mathrm{str}}$ with advantage $\varepsilon$ yields a PPT adversary $\mathcal{B}$ that wins the soundness game against $\Pi_{\mathrm{std}}$ with the same advantage $\varepsilon$ (up to negligible loss).
Proof.
We give a black-box reduction. $\mathcal{B}$ runs $\mathcal{A}$ while internally simulating the prover’s computations exactly as in $\Pi_{\mathrm{str}}$. By Theorem 4, the distribution of all commitments, openings, and auxiliary values output by $\Pi_{\mathrm{str}}$ is identical to those of $\Pi_{\mathrm{std}}$ on the same inputs and randomness. Therefore any accepting transcript $(x,\pi)$ produced by $\mathcal{A}$ against $\Pi_{\mathrm{str}}$ is an accepting transcript against $\Pi_{\mathrm{std}}$. Hence $\Pr[\mathcal{B}\ \text{wins against}\ \Pi_{\mathrm{std}}]=\Pr[\mathcal{A}\ \text{wins against}\ \Pi_{\mathrm{str}}]=\varepsilon$.
If, moreover, $\Pi_{\mathrm{std}}$ reduces soundness to the PCS binding property (as in standard analyses), then $\mathcal{B}$ composes with that reduction to yield a PCS-binding adversary with essentially the same advantage. No additional loss (such as a factor $\mathrm{poly}(T)$) is introduced by streaming/tree evaluation. ∎
Remark.
It is not sufficient to assume PCS binding alone implies global soundness; standard AIR/Plonkish systems rely on both PCS binding and algebraic soundness of randomized checks. The theorem above preserves whatever soundness guarantee the base protocol provides.
6.3 Zero-Knowledge
We adopt the standard indistinguishability game for (non-interactive) zero-knowledge.
Definition 12 (Zero-Knowledge Game).
$\mathsf{Game}_{\mathcal{A}}^{\mathrm{ZK}}(1^{\lambda},C)$:
  1. 1.
    The challenger generates $pp\leftarrow\mathsf{Setup}(1^{\lambda},D)$ and gives $pp$ to $\mathcal{A}$.
  2. 2.
    $\mathcal{A}$ outputs $(x,w)$ with $(x,w)\in R$.
  3. 3.
    The challenger samples $b\leftarrow\{0,1\}$.
  4. 4.
    If b=0b{=}0: $\pi\leftarrow\mathsf{Prover}_{\mathrm{str}}(pp,x,w)$.
  5. 5.
    If b=1b{=}1: $\pi\leftarrow\mathsf{Sim}_{\mathrm{str}}(pp,x)$.
  6. 6.
    $\mathcal{A}$ outputs $b^{\prime}$, and wins if $b^{\prime}=b$.
Theorem 9 (Zero-Knowledge Preservation).
Suppose the underlying protocol $\Pi_{\mathrm{std}}$ is computational zero-knowledge (e.g., honest-verifier ZK for the IOP compiled via Fiat–Shamir, using a hiding PCS and standard masking). Then $\Pi_{\mathrm{str}}$ is computational zero-knowledge with the same advantage bounds.
Proof.
Set $\mathsf{Sim}_{\mathrm{str}}:=\mathsf{Sim}_{\mathrm{std}}$. By Theorem 4, for any fixed public randomness (or programmed random oracle), the transcript distribution (commitments/openings and auxiliary values) of $\mathsf{Prover}_{\mathrm{str}}$ matches that of $\mathsf{Prover}_{\mathrm{std}}$; in non-hiding PCSs this equality is bit-for-bit. Consequently, the Fiat–Shamir challenge inputs have the same distribution, and any distinguisher against $\Pi_{\mathrm{str}}$ yields one against $\Pi_{\mathrm{std}}$ with the same advantage.
Instantiation note. If the PCS is not hiding by default (e.g., plain KZG), we assume a hiding instantiation or the standard polynomial-masking steps of the base protocol that render the overall scheme zero-knowledge. This transformation does not alter those steps; it only reorders computation and compresses memory. ∎
7 Performance and Implications
The sublinear-space prover does more than resolve a theoretical question; it alters the landscape of practical verifiable computation. In this section, we analyze its performance against existing systems and explore just a few new applications it enables.
7.1 Performance Comparison
The primary benefit of this construction is a square-root-space improvement: from linear space $\Theta(T)$ down to $O(\sqrt{T})$ (Section 5), i.e., a factor of $\approx\sqrt{T}$ less memory. [^1] Throughout, lower-order polylogarithmic terms are suppressed; see Theorem 5. To illustrate the practical impact, we compare the streaming prover against a standard linear-space prover for a ZKP system that materializes the full execution trace in memory.
Table 1: Performance Comparison: Linear-Space vs. Sublinear-Space Prover
Metric Standard Linear-Space Prover The Sublinear-Space Prover
Prover Space $\Theta(T)$ $O(\sqrt{T})$ (up to $\log\log$)
Prover Time $O(T\log N)$ or $O(T\log T)\dagger$ $O(T\log N)\cdot\mathrm{poly}(\lambda)$
[…]
Setup Model System-dependent (e.g., CRS/SRS) System-dependent (identical)
$\dagger$ Prover time is $O(T\log N)$ for KZG/IPA, or $O(T\log T)$ for STARK/FRI.
$\ddagger$ The “identical” entries hold for linear PCS instantiations (e.g., KZG/IPA), by Theorem 4. For hash-based STARK/FRI commitments, the prover remains sublinear-space but transcripts (and thus some constants) generally differ.
This construction dramatically improves the prover’s memory footprint while keeping time complexity comparable up to polylogarithmic factors. Crucially, the proof itself and the verifier’s work remain unchanged for linear PCS instantiations, ensuring that the prover can be a drop-in replacement in existing ZKP ecosystems without affecting end-users or protocol security parameters.
7.2 Concrete Impact on System Requirements
To contextualize the magnitude of the space improvement, consider two representative computation sizes. We assume each trace element occupies 32 bytes (e.g., a 256-bit field element), and suppress constant factors from the number of registers $k=O(1)$ and small auxiliary state. With $\lambda=\Theta(\log T)$, the dominant space bound is $O(\sqrt{T})$ up to a $\log\log$ factor (Section 5).
Scenario 1: Moderately large computation ($T=2^{24}\approx 1.68\times 10^{7}$)
  • Linear-space prover: $2^{24}\!\times 32=536{,}870{,}912$ bytes $\approx\mathbf{537}$ MB (decimal).
  • Streaming prover: $\sqrt{2^{24}}\!\times 32=2^{12}\!\times 32=131{,}072$ bytes $\approx\mathbf{0.13}$ MB.
Scenario 2: Very large computation ($T=2^{30}\approx 1.07\times 10^{9}$)
  • Linear-space prover: $2^{30}\!\times 32=34{,}359{,}738{,}368$ bytes $\approx\mathbf{34.4}$ GB (decimal).
  • Streaming prover: $\sqrt{2^{30}}\!\times 32=2^{15}\!\times 32=1{,}048{,}576$ bytes $\approx\mathbf{1.0}$ MB.
These examples highlight how this work makes verifiable computation accessible in scenarios where it was previously infeasible, moving from a specialized, server-bound task to one that can be performed on everyday devices.
7.3 Enabling New Deployment Scenarios
The removal of the linear-space barrier opens up several new frontiers for ZKP applications:
  • On-Device Proving: Mobile phones, IoT devices, and embedded systems can now generate proofs for complex computations, enabling applications like privacy-preserving health monitoring, on-device ML inference verification, and secure authentication without powerful backend servers.
  • Democratizing Trust: Lowering the hardware barrier allows a wider range of participants to generate proofs in decentralized networks (e.g., as provers in a ZK-rollup), enhancing security and decentralization.
  • Verifying Extremely Large Computations: Scientific simulations, large-scale data-processing pipelines, and complex financial models can now be made verifiable, as memory requirements no longer scale linearly with runtime.
7.4 Limitations and Future Work
This work presents both a significant advance and several exciting avenues for future research.
  • Tightening Space Bounds: The bound is $O(\sqrt{T}\cdot\log\log T)$ space (equivalently, $O(\sqrt{T})$ up to a $\log\log$ factor). It is natural to ask whether constants and lower-order terms can be further reduced, or whether matching lower bounds can be proved under general read-degree $r_{\text{reg}}$ and register count $k$.
  • Time Complexity Optimization: While the polylogarithmic overhead is acceptable in many memory-bound scenarios, reducing runtime remains valuable. Custom tree-evaluation routines tailored to AIR/PCS structure (e.g., avoiding generic interpolation where possible) could lower the $\log$ factors.
  • Hash-/lattice-based PCSs (STARK/FRI) and PQ settings: The identical-transcript guarantee relies on linear PCS commits (KZG/IPA). With Merkle/FRI-style hash commitments one can still stream with $o(T)$ space by carrying Merkle paths/FRI layers in $\mathsf{aux}$, but the resulting transcripts generally differ (different leaf chunking and digests). This raises $b_{\text{val}}$ (and thus the $\Delta\!\cdot\!b_{\text{val}}$ term in Cook–Mertz space). Quantifying the exact overhead while retaining sublinear space, and characterizing when transcript equivalence can be restored, are important open problems for post-quantum instantiations.
  • Beyond Proof Generation: The principle of decomposing cryptographic tasks into space-efficient tree evaluations may extend to other primitives (e.g., VDFs, interactive proofs), suggesting a broader algorithmic toolkit.
8 Conclusion
We have presented, to our knowledge, the first zero-knowledge proof system with sublinear prover space, breaking a long-standing barrier in verifiable computation. The core technical contribution behind this work is the Tree Evaluation Equivalence (Theorem 4), which establishes a mathematical equivalence between cryptographic proof generation and the Tree Evaluation problem. By combining the structural decomposition of Williams with a PCS whose commit map is linear, we express ZKP commitment generation as a recursive function $F_{v}$ over a tree, then evaluate it in low space using the Cook–Mertz algorithm. This yields a space improvement from $\Theta(T)$ to $O(\sqrt{T})$ up to a $\log\log$ factor.
The key algorithmic innovation is the commitment tree function $F_{v}$ (Algorithm 3), which composes cryptographic commitments while maintaining consistency via auxiliary channels. This construction replaces monolithic trace materialization with a space-efficient recursive evaluation that never stores the complete execution trace.
This construction preserves the security properties, proof size, and verification cost of existing linear-space systems, allowing it to serve as a direct, space-efficient replacement in a wide variety of applications (for linear PCS instantiations). Beyond the immediate practical impact, this work suggests a broader pattern of algorithmic transference, wherein advances in space-efficient algorithms for abstract computational models can be systematically lifted to create space-efficient cryptographic protocols for those same models. This observation points to a rich landscape where breakthroughs in algorithm design directly fuel improvements in cryptographic protocol efficiency and practical deployment.
The ability to generate proofs in sublinear space makes verifiable computation practical on resource-constrained devices and for computations of a scale previously thought infeasible. As our digital world increasingly demands strong guarantees of privacy and integrity, this work can provide a critical tool for making these guarantees accessible, affordable, and ubiquitous, while building a lasting bridge between computational complexity theory and applied cryptography.
Acknowledgements. We gratefully acknowledge Williams, and Cook &\& Mertz for their landmark results on square-root-space time–space simulations and the tree-evaluation perspective; this manuscript builds directly on these insights and would not exist without them. We further disclose that the exploration, drafting, and revision of this manuscript were conducted with the assistance of large language models (LLMs) for wording, copyediting, and organization; the authors bear sole responsibility for any errors in technical claims, constructions, and proofs. The authors declare that they have no conflicts of interest and received no external funding for this work.